Session: Feb 2001 (session 1)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (g.maccaferri at Wed Feb 14 09:21:14 MET 2001) Success

Torun (K.Borkowski at Wed Feb 14 10:25:27 MET 2001) Success
No known problems.

Effelsberg (A. Kraus at Wed Feb 14 11:06:06 MET 2001) Success

Medicina (g.maccaferri at Wed Feb 14 14:00:17 MET 2001) Success

Noto (C.Trigilio at Thu Feb 15 18:51:06 MET 2001) Success

Westerbork (Tony Foley at Fri Feb 16 16:34:41 MET 2001) Success
no known problems

Onsala (A. Polatidis at Wed Feb 21 01:50:22 MET 2001) Success

Jodrell_Bank (Alastair Gunn at Wed Feb 21 16:06:54 MET 2001) Minor failures
LO value set wrong between about 05:30 UT and 08:08 UT due to problems following FS crash. Data missing between these times.

Cambridge (Alastair Gunn at Wed Feb 21 16:08:25 MET 2001) Minor failures
Several motor fuses blew in VLBA recorder sometime between 03:45 UT and 05:22 UT. Recorder was not fixed before the end of the experiment. Data should be good until at least 03:45 UT; exact time of motor failure is not known.